MODS1 AGw Commissioning at LDG: 18 November 2010 UTC

Observer(s): R. Pogge and B. Shappee (OSU), M. Pedani and M. Wagner (LBTO)
Telescope Operator: G. Bechetti (LBTO) with D. Gonzalez in Tucson
Active Optics Support: J. Hill (LBTO) and D. Miller (LBTO remote from Tucson)
Telescope Support: J. Little
Instrument Support: D. Officer (LBTO)

Summary

2330 - Weather at handoff is beautifully clear, winds calm(!!!), Humidity 4%, Temp 4.5C, pressure rising. Looks like the promised high pressure system is here (sure blew in last night).

Conditions are photometric

This was the night we had been waiting for. Some of the best 30s science images were 0.52-arcsec FWHM in the SDSS r filter.

Acquired the AGw stage to PCS coordinate transformation data, field aberration measurements, and got preliminary 5-dice runs, astrometric and photometric data, and got to exercise the instrument heavily.

The telescope was well-behaved. We lost only ~15 minutes to a problem with the LDG rotator struggling to keep up at EL=88.

In all, a great night.

Preparations

2330 - MODS dewars filled - still showing excellent hold time

Issues with the IDL scripts appear to be understood (FITS header info in the WFSC images is at fault). Doug Miller is making modifications for a workaround. See Issue 3011.

Details

2330 - Starting imaging and spectral mode bias frames

3Kx3K Bias: mods1x.20101118.0001-0010

8Kx3K Bias: mods1x.20101118.0011-0020

Startup

0020 - As twilight approaches, Geno's starting the telescope without too much trouble. Conditions are quite good at sunset.

MODS startup procedure for telescope start-of-observing setup
  • Open the Hatch
  • Retract the Calibration Tower
  • Reset (home) the AGw Stage
  • Send the AGw Stage to field center

This will be automated by a script and become part of the suite of tools.

0045 - Pointing Check

Coefficients:
    IE = 23
    CA = 23

0050 - Collimation Setup

BS9183 - doing adjustment by hand so far.

WFS correction,m seeing is 1arcsec but we're not yet equilibrated. Much more stable seeing that previous nights.

Conditions are photometric, but warming up outside, now about 3.4C, humidity 6%, slowly rising pressure, wind 4m/s and steady out of the East

Photometric Calibration

0100 - Going to the M71 SDSS calibration field while waiting for the telescope to settle to better seeing for the transformations

Seeing about 1arcsec. On the star, the GCS started out not being able to center it. I got better performance by putting in the F525LP filter. Is it a background level problem? Not obviously the cause, but just sayin...

Centering IMCS up, looks good and quick lock-on (30sec).

Taking 10s snapshot in g and r to gauge exptime.

Looks good, running photcal.pro to get ugriz snaps:
  • 2x(g+r) 30s and 10s
  • u + 2x(i) 60 and 30s
  • u + 2x(z) 60 and 30s

Images: mods1s.20101117.0025 to 32 in red, 30 in blue

0127 - Some images g and r have 0.7 and 0.8-arcsec round images. finally...

The images are lovely. smile

The SDSS z images are 0.6-arcsec FWHM and round. And we still haven't yet co-focused the instrument and telescope! I'm not sure I want to try. Sweet.

AGw Stage Transformation - this time for-real??

The wfs images are in /OldRepository/AGW_Data/20101118/

Stone_A1

0135 - Preset to Stone A1, just to collimate while John prepares the script...

0150 - Running transform_collect, imcs set, 10s integrations in sdss r in the science camera.

0159 - Star selection had a few outside the guide patrol field, so script failed. John is recasting the star list, using the LBTView tool as a guide.

Lots of moonlight scatter in the images, the telescope webcam shows a huge glint. We're not that far from the Moon.

0212-0220 - John having problems with GCS, getting RPC errors of some kind.

0227 - only 3 stars at PA=-90 due to moon and WFS problems

Star 1 at PA=-90 is OK, stars 4 and 7 are poorly guided due to moonlight background (still with LP filter)

0229 - Running PA=0

Star 6 was completely swamped by moonlight.

Seeing is now 1.1 to 1.2 arcsec on this field, El=48. Guider is reporting 1.1 - 1.4.

Star 7, 8 was mis-acquired from a moon glint.

Stone_O

0249 MODS preset to Stone O

We called Doug to twist his arm to give us a StoneO catalog that is farther from the moon. Stop and restart IDL to load the new catalog.

10-sec MODS image has 0.9 arcsec FWHM in r.

02:58 Clear Active Optics on M1 - this removes some of the residual spherical that had accumulated. Doug was looking at the pupil positioning. And we note that the temperature is rising at a rate of 1 degC/hour.

03:09 transform_collect, 'Stone_O.list', /MODS, START=1
starting with wfsc000028 and mods1r.20101118.0049

Star 24 probably acquired a neighboring star.
Couldn't move probe to star 22 (was a typo - should have been 23). Star 23 probably acquired a neighboring star.

LBTO@obs2:3 % cat Stone_O.list
Stone_O
1   0
11  0
15  0
24  0
23  0
28  0
26  0
33  0
32  0
12  0
10  0
1   10
11  10
15  10
24  10
23  10
28  10
26  10
33  10
32  10
12  10
10  10

0430 Finished 11 stars at each of three PAs for transformation data. Last wfsc000086 and mods1r.20101118.0081.

Photometry - high airmass shot

0433 - Back to M71 Field B. At airmass 2.85 (we wanted high extinction...)

Re-executing photcal.pro from earlier...

Images: mods1b.20101118.0031 and mods1r.20101118.0082 et seq.

0445 - Last images blue: 0036, red: 0089

Field Aberration Measurement

0450 - Pointing to BS9107, on-axis active preset

The goal in target selection is to avoid Mr. Moon at 0h/+5d.

Temperature has been rising, now 6C (!), Wind 6 m/s steady from SE, Mirror at +3.6C, Humidity is 4%, still slow rise in pressure.

Interesting night...

Seeing is 0.8 arcsec on GCS. Nice...

0505 - Shot images of the field to see the guide probe obscuration on-axis while waiting on the script prepation by JH.

Field Aberration Data on BS9107 - almost

John creates a new field star list for BS9107 to match the new star numbers in the new catalog.

05:13 field_collect, 'BS9107_field.list'

But this fails because the script has wrong Y values for the stars (wrong position angle). The preset for star 23 was sent at PA=28.83 when is should have been 208.83 (or -151.2).

Doug on Skype fixes the problem quickly. The correct position angle for MODS was being calculated, but it wasn't being sent in the preset.

Data files live (along with the transformation data in /OldRepository/AGW_Data/20101118/

0549 field_collect, 'BS9107_field.list'

first on-axis image is left_wfsc000097

Seeing on the guider is bouncing between 0.7 and 1.2 arcsec.

Acquisition of star 50 failed because guiding not stable. Why? Y was too large, so the probe offset failed.

0608 field_collect, 'BS9107_field.list', START=4

This group starts with left_wfsc000111

Seeing has improved to 0.6 to 0.8 on the guider.

Pointing has drifted with temperature, so we send an acquire preset (from Geno's IRAF session). IE=+7 CA=+12.

0623 field_collect, 'BS9107_field.list', START=5

starting with left_wfsc000117

M2Total X=-4.01 Y=2.683 Z=0.0 RX=148.65 RY=181.8
M1Total X=1.034 Y=1.904 Z=0.226 RX=-23.66 RY=2.49 at 82.7
M1Global X=0 Y=+3.00 Z=-0.1 RX=0 RY=0
M1 and M2 were parallel shifted relative to the collimation model to compensate the rising temperature.

GS 45 missed the probe offset so guiding wasn't stable.

0642 field_collect, 'BS9107_field.list', START=9

This group starts with left_wfsc_000132

Rick reports 0.53 arcsec round images in 30-sec r' and 0.6 in g' on MODS. Now he's getting anxious to take science data. Guider is still reporting 0.7, so we should figure out why it is so pessimistic.

0715 Acquire preset to adjust pointing, but we kept it the same.

Update the list of stars to remove 45, 50.

BS9107 Field Data - second pass

0718 field_collect, 'BS9107_field.list', START=1

For this group, the first image is left_wfsc000161

Seeing is 0.6 to 0.8 on the guider. 0.61 arcsec in g', 0.58 in r' on MODS.

Outside temperature is down to +3.2 degC after a peak of 6 degC.

0818 Star 47 hits the Y-limit and preset fails for unstable guiding. This causes only a momentary (1 min) disruption in the data collection.

0819 field_collect, 'BS9107_field.list', START=12

This group starts at left_wfsc000214

The list with the bad stars weeded out is: BS9107_field.list
BS9107
23
28
36
33
42
39
43
44
41
22
24

left_wfsc000219 looks to be saturated. Why?

The last on-axis image was left_wfsc000225.

Andrew's Summary of the Field Aberration Data
> From: Andrew Rakich 
> Date: November 18, 2010 5:05:11 PM MST
> Subject: Re: MODS1 AGw transformation and field aberration data
>
> Well, my preliminary conclusion is that we have focal plane tilt without any obvious binodal astigmatism.
>
> I need to look at the nuts and bolts of the data more, but at this stage I would not reccommend adjusting M2 at all.

Focus Tests

0831 MODS preset on BS9107 for camera focus test

Running red camera 1150 to 1350 in steps of 50 microns, blue camera left at nominal focus as the fixed reference. 10s exposures

Files: mods1r.xxx.0098, mods1b.xxx.0045

Best focus is 1250-microns, the nominal focus we had before, so we nailed it good in the lab!

Best seeing was 0.52-arcsec FWHM.

Five-Dice offset runs

0845 - Running the Five-Dice sequences since seeing is good and we have time before the Stone E transform field comes into play. We should be able to run a low- and high-elevation set before getting back to transformation measurments.

0852 - games with azimuth wanting 448-degrees on a dec=+31 preset. We instead are going to our low elevation field C0631-19 (Dec=-19), target elevation 35-deg

0856 - Primary hit the limit, subtracted a 1mm global X offset and resending the preset after it is done

0900 - Back to stable, starting dice10.pro (5x 5-dice pattern with 10-arcsec throw)

Seeing at this elevation is ~0.8 arcsec

Images: mods1r.20101118.0103 - 0127

0927 - C0630+31, El=83 degrees. Seeing is

Started dice10.pro, images: mods1r.20101118.0128 - 153

0959 - Telescope offset command stalled and timed out after 300s. Something goofy about the rotator. Geno & John resetting the telescope to try to recover.

We lost the last 5-dice sequence of 5, but we're OK for the high/low rehearsal. Last image was 150. JMH: Tom Sargent says the rotator started misbehaving at EL=85. This might have been the offsets settling, or those images might be funny.

The offset timed out after 5 minutes. Problem was apparently that rotator not being able to keep up at about El=88 degrees. The fault message was not very informative of what was really going on. Geno is opening IssueTrak #3010.

Transformation Measurement - Stone E field

1008 - IMCS setup and Active collimation

1013 - transform_collect, 'Stone_E.list', /MODS, START=1

First wfs image is left_wfsc000226 at EL=56. Seeing on guider is 0.8 - 1.2 arcsec FWHM. These images are at PA=-90.

1048 - transform_collect, 'Stone_Em80.list', /MODS, START=1

First wfs image is left_wfsc000248 at EL=57. These images are at PA=-80. First MODS image in the second set is mods1r.20101118.0169. Two of these stars at the new PA could not be reached by the guide probe.

Last wfs image is left_wfsc000270 Last MODS image is mods1r.2010.0184

AGw Guide Camera Astrometry

1140 - Taking guide camera images in the Orion Nebula core cluster with Jones and Walker (1988) astrometry. Will use to calibrate the pixel scale and other parameters of the guide camera CCD.

Much hunting about as we had the wrong coords in the table - used MODS science field and the Jones & Walker finder (Orion Core is too blasted to use the LBTView tool)

Shot 20 images, /Old/Repository/20101118_MODS/left_guider000001-21.fits, exptime=5s, F525LP filter.

M67 Photometric Field

1205 - Offset field in M67 08:50:55/11:52:15 away from the brightest stars which would badly saturate the science CCDs and obliterate stars. About 21 high-precision ugriz stars in this field.

1211 - Star is elliptical looking, but WFS looks fine. Clearing C00 on M1. This improves the "hair" on the guide star image, but it is still elliptical. Not that the WFS subap images are also elliptical - must be a binary star. See Issue 3013.

1216 - Not much help - might be a bad star (two stars?) Picking another guide star (8:50:47.69/11:47:52.74) This one is round.

1218 - running photcal.,pro, images: mods1b.xxx.0050 and mods1r.xxx.0189

1230 - Finished. Closing up.

-- RichardPogge - 17 Nov 2010
Topic revision: r14 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback